The price of a 2-hour plumbing repair job is $130 and the price of a 5-hour job is $250. Find and interpret the slope of the line through (2, 130) and (5, 250).
Answer by stanbon(75887)

The price of a 2-hour plumbing repair job is $130 and the price of a 5-hour job is $250. Find and interpret the slope of the line through (2, 130) and (5, 250).
-------------
Your "x" values are time.
Your "y" values are cost
------------------------------
slope = [change in cost/change in time] = (250-130)/(5-2) = 120/3 = 40
----------------------------------------
This slope tells you the cost increases by $40 for each additional
hour of work.
============================
Cheers,
Stan H.